Northwood Company manufactures basketballs. The company has a ball that sells for $36. At present, the ball is manufactured in a small plant that relies heavily on direct labor workers. Thus, variable expenses are high, totaling $21.60 per ball, of which 60% is direct labor cost.

    Last year, the company sold 56,000 of these balls, with the following results:
  Sales (56,000 balls) $ 2,016,000
  Variable expenses 1,209,600
  Contribution margin 806,400
  Fixed expenses 662,400
  Net operating income $ 144,000
Required:
1-a.

Compute last year's CM ratio and the break-even point in balls. (Do not round intermediate calculations.)

  

1-b.

Compute the the degree of operating leverage at last year’s sales level. (Round your answer to 2 decimal places.)

  

2.

Due to an increase in labor rates, the company estimates that next year's variable expenses will increase by $2.88 per ball. If this change takes place and the selling price per ball remains constant at $36.00, what will be next year's CM ratio and the break-even point in balls? (Do not round intermediate calculations.)

  

3.

Refer to the data in (2) above. If the expected change in variable expenses takes place, how many balls will have to be sold next year to earn the same net operating income, $144,000, as last year? (Do not round intermediate calculations. Round your answer to the nearest whole unit.)

  

4.

Refer again to the data in (2) above. The president feels that the company must raise the selling price of its basketballs. If Northwood Company wants to maintain the same CM ratio as last year (as computed in requirement 1a), what selling price per ball must it charge next year to cover the increased labor costs? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

  

5.

Refer to the original data. The company is discussing the construction of a new, automated manufacturing plant. The new plant would slash variable expenses per ball by 40%, but it would cause fixed expenses per year to increase by 80%. If the new plant is built, what would be the company’s new CM ratio and new break-even point in balls? (Do not round intermediate calculations.)

  

6.

Refer to the data in (5) above.


a.

If the new plant is built, how many balls will have to be sold next year to earn the same net operating income, $144,000, as last year? (Do not round intermediate calculations. Round up your final answer to the nearest whole number.)

b-1.

Assume the new plant is built and that next year the company manufactures and sells 56,000 balls (the same number as sold last year). Prepare a contribution format income statement (Do not round your intermediate calculations.)

b-2. Compute the degree of operating leverage. (Do not round intermediate calculations and round your final answer to 2 decimal places.)
Answer #1

1. a. CM ratio = 100% -60% = 40%

Break even point = Fixed expenses/ (Selling price - Variable cost) = 662400 / ( 36-21.6) = 46000 balls

b. DOL = Contribution margin/ Net operating income = (806,400/ 144,000)*100 = 5.6 times

2.a. CM ratio = Contribution / Selling price *100 = ( 36-24.48) / 36 * 100 = 32%

b. Break even point = Fixed expenses/ (Selling price - Variable cost) = 662400 / 11.52 = 57500 balls

3. No of balls to be sold = Fixed costs + Desired Income / Contribution per unit = (662400+ 144000) / 11.52 = 70000 balls

4. Let selling price be x

(x - 24.48) = 0.4x

x = 40.8

5. Variable cost = 21.6 * 60% = 12.96

Fixed costs = 662400 * 180% = 1192320

a. CM ratio = Contribution / Selling price *100 = ( 36-12.96) / 36 * 100 = 64%

b. Break even point = Fixed expenses/ (Selling price - Variable cost) = 1192320 / 23.04 = 51750 balls

6.

a. No of balls to be sold = Fixed costs + Desired Income / Contribution per unit = (1192320+ 144000) / 23.04 = 58000 balls

b1.

Sales (56,000 balls) 2016000
Variable expenses 725760
Contribution margin 1290240
Fixed expenses 1192320
Net operating income 97920

b2. DOL = Contribution margin/ Net operating income = (1290240/ 97920)*100 = 13.17 times